If a buffer solution is 0.420 M in a weak acid (Ka = 5.6 × 10-6) and 0.170 M in its conjugate base, what is the pH?
Answer #1

the buffer

pH = pKa + log(conjugate base / acid)

pKa = -log(KA) = -log(5.6*10^-6) =5.2518

pH = pKa + log(conjugate base / acid)

pH = 5.2518+ log(0.17/0.42) = 4.8589